OpenStudy (anonymous):

Nancy has a sewing project which calls for panels which use 1 1/4 yards of fabric each. She has 4 1/2 yd of fabric. How many panels can she cut from the fabric? Answer 2 3 4 5

OpenStudy (dumbsearch2):

= (4 1/2)/(1 1/4) = (9/2)/(5/4) = 9/2(4/5) or 36/10 = 18/5 or 3 3/5 Answer: 3 panels

OpenStudy (dumbsearch2):

Mark my answer as correct if it helped. ^_^ By clicking "best answer".